9.Findarelationbetweenxandyifthepoints(x,y),(1,2)and(7,0)arecollinear.
 

Grade:10

1 Answers

Aditya Gupta
4 years ago
since the 3 points lie on a line, their slopes must be equal.
let A= (x,y), B= (1,2) and C= (7,0).
slope of AB= (y – 2)/(x – 1) and slope of BC= (2 – 0)/(1 – 7)= – 1/3
so, equating the 2,
(y – 2)/(x – 1)= – 1/3
or 3y+x= 7
kindly approve :=)